OpenStudy (anonymous):

Sally has only nickels and dimes in her money box. She knows that she has less than $20 in the box. Let x represent the number of nickels in the box and y represent the number of dimes in the box. Which of the following statements best describes the steps to graph the solution to the inequality in x and y?

OpenStudy (anonymous):

Draw a dashed line to represent the graph of 5x + 10y = 2000, and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 10x + 5y = 2000, and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 5x + 10y = 2000, and shade the portion below the line for positive values of x and y. Draw a dashed line to represent the graph of 10x – 5y = 2000, and shade the portion below the line for positive values of x and y.

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

since 1 nickel= 5 cent, and 1 dime = 10 cent, then the total amount of money of Sally, is: \[5x + 10y\]

OpenStudy (anonymous):

Okay

OpenStudy (michele_laino):

and that amount of money has to be less than $20, namely less than 2000 cent

OpenStudy (anonymous):

okay, so would it be shaded below the line then?

OpenStudy (michele_laino):

yes!

OpenStudy (michele_laino):

since we have this inequality: \[5x + 10y < 2000\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

Wow you have explained it perfectly!

OpenStudy (michele_laino):

thanks! :)

OpenStudy (anonymous):

No, no, thank you!

OpenStudy (michele_laino):

:)